# Proximal Interacting Particle Langevin Algorithms

### v2

@@ A.4.2 Convergence and discretisation bounds
$\nabla G_{1}(z_{\theta},z_{1},\dots,z_{N})=$ $\Big{(}N^{-1}\sum_{i=1}^{N}\nabla_{\theta}g_{1}(z_{\theta},\sqrt{N}z_{i}),N^{-1/2}\nabla_{z_{1}}g_{1}(z_{\theta},\sqrt{N}z_{1}),\dots,$
$N^{-1/2}\nabla_{z_{N}}g_{1}(z_{\theta},\sqrt{N}z_{N})\Big{)}^{\intercal}$
@@
Theorem A.7 (Theorem 4.2 restated).
Let A1A5 and C1 hold. Then for $\gamma\leq 1/L_{g_{1}}$ and $\gamma\leq\lambda\leq\gamma/(1-\mu\gamma)$, PIPGLA satisfies
@@ Appendix B Theoretical Analysis of Proximal Particle Gradient Descent
B.1 Background on Particle Gradient Descent of Kuntz et al. [2023]
The PGD algorithm relies on the perspective that the MMLE problem can be solved by minimising the free energy
@@ Appendix C Convergence to Wasserstein gradient flow
We start by proving the following auxiliary result. Let us denote, for any $\theta\in^{d_{\theta}}$ and $\nu\in{\mathcal{P}}(^{d_{x}})$, $g(\theta,\nu):=\int_{{}^{d_{x}}}\nabla_{\theta}U^{\lambda}(\theta,x^{\prime})\nu(x^{\prime})\mathrm{d}x^{\prime}$.
@@
The function $g:^{d_{\theta}}\times{\mathcal{P}}(^{d_{x}})\to\mathbb{R}^{d_{\theta}}$ is Lipschitz continuous in both arguments, i.e.,
@@
Rockafellar and Wets [2009, Proposition 12.19] shows that $\nabla U^{\lambda}$ is Lipschitz continuous with Lipschitz constant $\lambda^{-1}$. Then the result follows from Akyildiz et al. [2023, Lemma 5]. ∎
@@
The proof exploits the Lipschitz continuity of $\nabla U^{\lambda}$ and of $g$ established in Lemma C.1. The argument is classical and omitted, see Akyildiz et al. [2023, Proposition 8] for the proof in a similar context. ∎
@@
from which the result follows.
Appendix D Experiments Details
D.1 Derivation of the proximal operators
D.1.1 Laplace prior with unknown mean $\theta$
@@
We recall that $\phi\in^{d}$ is a subdifferential of the $\ell^{1}$-norm at $x\in^{d}$ if and only if $\phi_{i}(x)=\text{sign}(x_{i})$ if $x_{i}\neq 0$ and $|\phi_{i}(x)|\leq 1$ otherwise [Parikh and Boyd, 2014].
@@
D.1.2 Laplace prior with unknown scale $e^{2\theta}$
For the Bayesian neural network experiment we consider a Laplace prior with zero mean and unknown scale parameterised by $e^{2\theta}$ (which ensures that the scale is positive), we have $g_{2}(\theta,x)=2d_{x}\alpha+\sum_{i}|x_{i}|e^{-2\alpha}$. Its proximal operator is given by
$\operatorname{prox}_{g_{2}}^{\lambda}(\theta,x)=\operatorname*{arg\,min}_{(u_{0},u)}h(u_{0},u),\quad h(u_{0},u)=2u_{0}d_{x}+\sum_{i}|u_{i}|e^{-2u_{0}}+\|(u_{0},u)-(\theta,x)\|^{2}/{(2\lambda)}.$
@@ D.1.2 Laplace prior with unknown scale $e^{2\theta}$
$0\in\partial\big{(}2u_{0}d_{x}+\sum_{i}|u_{i}|e^{-2u_{0}}\big{)}+\nabla\big{(}\|(u_{0},u)-(\theta,x)\|^{2}/(2\lambda)\big{)},$
@@
00 0 $=2d_{x}-2e^{-2u_{0}}\sum_{i=1}^{d_{x}}|u_{i}|+\frac{1}{\lambda}(u_{0}-\theta),$
@@
$u_{0}$ $=\theta-2\lambda d_{x}+2\lambda e^{-2u_{0}}\sum_{i}|u_{i}|$ (53)
@@
$u_{0}\approx\theta-2\lambda d_{x}+\frac{1}{2}W(4\lambda e^{-2\theta}\sum_{i}|u_{i}|).$
D.1.3 Uniform prior
@@
D.1.4 Approximation for PIPULA and PPGD
@@
D.2 Bayesian logistic regression
Dataset
We create a synthetic dataset by first fixing the value of $\theta$ and sampling the latent variable $x\in\mathbb{R}^{50}$ from the corresponding prior. We then sample the 900 observations from a Bernoulli distribution with parameter $s(v_{j}^{T}x)$, where $s$ is the logistic function and the entries of the covariates $v_{j}$ are drawn from a uniform distribution $\mathcal{U}(-1,1)$. The true value of $\theta$ is set to $\theta=-4$ for the Laplace prior and $\theta=1.5$ for the uniform one.
@@ D.2 Bayesian logistic regression
Results
Figure 3 shows the $\theta$-iterates obtained with MYIPLA and PIPGLA starting from 7 different initial values $\theta_{0}$ and using the approximate solver for $\operatorname{prox}_{g_{2}}^{\lambda}$ with $g_{2}(\theta,x)=\sum_{i=1}^{d_{x}}|x_{i}-\theta|$ and an iterative procedure using 40 iterations in each step. We observe that the iterative solver results in a slightly slower convergence to stationarity, but overall the two sets of algorithms converge to the same true value of $\theta$. We also observe that the convergence to stationarity for PIPGLA is much slower compared to MYIPLA. However, if we increase the value of $\gamma$ in the hope of faster convergence, the iterates either do not converge to the true value or the standard deviation is significantly larger. For all algorithms considered, approximate solvers are 25% faster than iterative solvers (see Table 2).
We also compare the results for the Uniform prior, in this case we only use the approximate proximity map (Figure 4).
@@ Results
(c) PIPGLA $(\lambda=\gamma)$, approximated $\operatorname{prox}_{g_{2}}^{\lambda}$.
(d) PIPGLA $(\lambda=\gamma)$, iterative solver for $\operatorname{prox}_{g_{2}}^{\lambda}$.
@@
Since all the algorithms considered aim at estimating the parameter $\theta$ by sampling from a distribution which concentrates around $\theta^{\star}$, we compare the estimators of $\theta^{\star}$ obtained by using only the last iterate $\theta_{K+1}^{N}$ and averaging over a number of iterates. We compare the normalised MSE (NMSE) for $\theta$ for the estimator obtained by averaging the $\theta$-iterates after discarding a burn-in of 1500 samples (column named avg) against using the last $\theta$ of the chain (column last). The results are in agreement, with the NMSE for the averaged estimator having lower variance in most settings (Table 8).
Table 8: Bayesian logistic regression for Laplace and uniform priors. Normalised MSE (NMSE) for the last iterate of $\theta$ (last) and the posterior mean after discarding a burn-in of 1500 samples (avg). Each different algorithm is run 500 times for different starting points using 50 particles and 5000. NMSEs are averaged over the 500 replicates. The second column indicates whether the proximal map is calculated approximately or iteratively, using 40 steps in each iteration. For the uniform prior case we have not implemented the iterative method.
@@
D.3 Bayesian neural network
@@
We use the MNIST dataset. Features are normalised so that each pixel has mean zero and unit standard deviation across the dataset. We split the dataset into 80/20 training and testing sets.
@@
$\operatorname{prox}_{g_{2}}^{\lambda}(\alpha,w)=\operatorname*{arg\,min}_{(u_{0},u)}h(u_{0},u),\quad h(u_{0},u)=2u_{0}d_{w}+\sum_{i}|u_{i}|e^{-2u_{0}}+\|(u_{0},u)-(\alpha,w)\|^{2}/{(2\lambda)},$
whose approximate solution is calculated in Section D.1.2.
@@
For the $x$-gradients of $g_{1}$, we use JAX’s grad function (implementing a version of autograd). Plugging the expressions above in the corresponding equations, we can implement the proposed algorithms. However, due to the high dimensionality of the latent variables, we stabilise the algorithm using the heuristics discussed in Section 2 of Kuntz et al. [2023]. This simply entails dividing the gradients and proximal mapping terms of the updates of $\alpha$ and $\beta$ by $d_{w}$ and $d_{v}$. We then set $\gamma=0.05$ and $\lambda=0.5$ which ensures that the algorithms are not close to losing stability. In addition, the weights of the network are initialised according to the assumed prior. This is done by setting each weight to ±$\pm\;a\log u$ where $u\sim\mathcal{U}(0,1)$, the sign is chosen uniformly at random and $a>0$ is interpreted as the average initial size of the weights. Williams [1995] suggests setting $a=1/\sqrt{2m}$ for $w$ and $a=1.6/\sqrt{2m}$ for $v$, where $m$ is the fan-in of the destination unit.
@@
and the log pointwise predictive density (LPD, Vehtari et al. [2017])
$\text{LPD}:=\frac{1}{|\mathcal{Y}_{\text{test}}|}\sum_{(f,l)\in\mathcal{Y}_{\text{test}}}\log(g(l|f)).$
Under the assumption that data is drawn independently from $p(\mathrm{d}l,\mathrm{d}f)$, we have the following approximation for large test data sets,
LPD $\approx\int\log(g(l|f))p(\mathrm{d}l,\mathrm{d}f)=\int\bigg{[}\int\log\Big{(}\frac{g(l|f)}{p(l|f)}\Big{)}p(\mathrm{d}l|\mathrm{d}f)\bigg{]}p(\mathrm{d}f)+\int\log(p(l|f))p(\mathrm{d}l,\mathrm{d}f)$
@@
This means that the larger the LPD is, the smaller the mean KL divergence between our classifier $g(l|f)$ and the optimal classifier $p(l|f)$.

### v3

@@ A.4.2 Convergence and discretisation bounds
$\nabla G_{1}(z_{\theta},z_{1},\dots,z_{N})=\left(N^{-1}\sum_{i=1}^{N}\nabla_{\theta}g_{1}(z_{\theta},\sqrt{N}z_{i}),N^{-1/2}\nabla_{z_{1}}g_{1}(z_{\theta},\sqrt{N}z_{1}),\dots,N^{-1/2}\nabla_{z_{N}}g_{1}(z_{\theta},\sqrt{N}z_{N})\right)^{\intercal}$
@@
Theorem A.15.
[Theorem 4.2 restated] Let A1, A2, A4 and C1 hold. Then for $\gamma\leq 1/L_{g_{1}}$ and $\gamma\leq\lambda\leq\gamma/(1-\mu\gamma)$, PIPGLA satisfies
@@ Appendix B Theoretical Analysis of Proximal Particle Gradient Descent
B.1 Background on Particle Gradient Descent
The PGD algorithm [Kuntz et al., 2023] relies on the perspective that the MMLE problem can be solved by minimising the free energy
@@ Appendix C Convergence to Wasserstein gradient flow
We start by proving the following auxiliary result. Let us denote, for any $\theta\in\mathbb{R}^{d_{\theta}}$ and $\nu\in{\mathcal{P}}(\mathbb{R}^{d_{x}})$, $g(\theta,\nu):=\int_{\mathbb{R}^{d_{x}}}\nabla_{\theta}U^{\lambda}(\theta,x^{\prime})\nu(x^{\prime})\mathrm{d}x^{\prime}$.
@@
The function $g:\mathbb{R}^{d_{\theta}}\times{\mathcal{P}}(\mathbb{R}^{d_{x}})\to\mathbb{R}^{d_{\theta}}$ is Lipschitz continuous in both arguments, i.e.,
@@
Rockafellar and Wets [2009, Proposition 12.19] shows that $\nabla U^{\lambda}$ is Lipschitz continuous with Lipschitz constant $\lambda^{-1}$. Then the result follows from Akyildiz et al. [2025, Lemma 5]. ∎
@@
The proof exploits the Lipschitz continuity of $\nabla U^{\lambda}$ and of $g$ established in Lemma C.1. The argument is classical and omitted, see Akyildiz et al. [2025, Proposition 8] for the proof in a similar context. ∎
@@
from which the result follows. ∎
Appendix D Algorithm Comparison
We provide further details on computing the complexity estimates of Section 4.5. For convenience, we summarise the complexity estimates for $\lambda$, the number of particles $N$, $\gamma$, and the number of steps $n$ to achieve an error $\mathbb{E}\left[\|\theta_{n}^{N}-\bar{\theta}^{\star}\|^{2}\right]^{1/2}=\mathcal{O}(\varepsilon)$ from Table 1. The values are provided in terms of the key parameters $d_{\theta},d_{x}$ and $\delta>0$ is any small positive constant.
$\lambda$ $N$ $\gamma$ $n$
On the other hand, we also compared the algorithms in terms of their computational requirements. We account for the computational cost of running each algorithm for $n$ iterations with $N$ particles and time discretisation step $\gamma$, while guaranteeing an $\mathcal{O}(\varepsilon)$ error, in terms of component-wise evaluations of $\nabla g_{1}$ and $\operatorname{prox}_{g_{2}}^{\lambda}$, and number of independent standard $1$-dimensional Gaussian samples.
Finally, Table 4 summarises the key differences and advantages of each algorithm. We recall that $L_{g_{1}}$ and $\mu$ denote the Lipschitz continuity and strong-convexity parameters of $g_{1}$, introduced in assumptions A1 and A4, respectively.
[…]
Table 4: Comparison of convergence assumptions, parameter constraints, and advantages of each algorithm.
Assumptions for Constraints on $\lambda$ Constraints on $\gamma$ Advantages
convergence
[…]
helps escape local minima.
PIPGLA A1, A2, A4 and C1 $\gamma\leq\lambda\leq\gamma/(1-\mu\gamma)$ $\gamma\leq 1/L_{g_{1}}$ Ensures $\theta$ estimates remain
within the support of the distribution.
MYPGD A1A5 $\lambda\geq 0$ $\gamma<(L_{g_{1}}+\lambda^{-1}+\mu)^{-1}$ Produces lower-variance estimates
for MMLE in the strongly convex setting.
Appendix E Numerical Experiments
E.1 Derivation of the proximal operators
E.1.1 Laplace Prior with Unknown Mean $\theta$
@@
We recall that $\phi\in\mathbb{R}^{d}$ is a subdifferential of the $\ell^{1}$-norm at $x\in\mathbb{R}^{d}$ if and only if $\phi_{i}(x)=\text{sign}(x_{i})$ if $x_{i}\neq 0$ and $|\phi_{i}(x)|\leq 1$ otherwise [Parikh and Boyd, 2014].
@@
E.1.2 Laplace Prior with Unknown Scale $e^{2\theta}$
For the Bayesian neural network experiment we consider a Laplace prior with zero mean and unknown scale parameterised by $e^{2\theta}$ (which ensures that the scale is positive), we have $g_{2}(\theta,x)=d_{x}\alpha+\sum_{i}|x_{i}|e^{-2\alpha}$. Its proximal operator is given by
$\operatorname{prox}_{g_{2}}^{\lambda}(\theta,x)=\operatorname*{arg\,min}_{(u_{0},u)}h(u_{0},u),\quad h(u_{0},u)=u_{0}d_{x}+\sum_{i}|u_{i}|e^{-2u_{0}}+\|(u_{0},u)-(\theta,x)\|^{2}/{(2\lambda)}.$
@@ E.1.2 Laplace Prior with Unknown Scale $e^{2\theta}$
$0\in\partial\big{(}u_{0}d_{x}+\sum_{i}|u_{i}|e^{-2u_{0}}\big{)}+\nabla\big{(}\|(u_{0},u)-(\theta,x)\|^{2}/(2\lambda)\big{)},$
@@
00 0 $=d_{x}-2e^{-2u_{0}}\sum_{i=1}^{d_{x}}|u_{i}|+\frac{1}{\lambda}(u_{0}-\theta),$
@@
$u_{0}$ $=\theta-\lambda d_{x}+2\lambda e^{-2u_{0}}\sum_{i}|u_{i}|$ (50)
@@
$u_{0}\approx\theta-\lambda d_{x}+\frac{1}{2}W\left(4\lambda e^{-2\theta}\sum_{i}|u_{i}|\right).$
E.1.3 Uniform Prior
@@
E.1.4 Approximation for PIPULA and PPGD
@@
E.2 Bayesian logistic regression
and for the uniform prior, we obtain
where $g_{1}$ is differentiable and $g_{2}$ is lower semi-continuous, and $\imath_{\mathcal{K}}$ is the convex indicator of $\mathcal{K}$ defined by $\imath_{\mathcal{K}}(x)=0$ if $x\in\mathcal{K}$ and $\imath_{\mathcal{K}}(x)=\infty$ otherwise.
In both cases we have that
where $C$ is a constant. As shown in Akyildiz et al. [2025, Section 6.1.1], the function $g_{1}$ is gradient Lipschitz and strictly convex but not strongly convex. The function $g_{2}$ satisfies A1 for both the Laplace and the uniform prior, as observed in Pereyra [2016], in the case of the Laplace prior $g_{2}$ also satisfies A3 while the uniform prior does not lead to a Lipschitz $g_{2}$. Since $g_{1}$ does not depend on $\theta$, A5 holds for the Laplace prior.
Dataset.
We create a synthetic dataset by first fixing the value of $\theta$ and sampling the latent variable $x\in\mathbb{R}^{50}$ from the corresponding prior. We then sample the 900 observations from a Bernoulli distribution with parameter $s(v_{j}^{T}x)$, where $s$ is the logistic function and the entries of the covariates $v_{j}$ are drawn from a uniform distribution $\mathcal{U}(-1,1)$. The true value of $\theta$ is set to $\bar{\theta}_{\star}=-4$ for the Laplace prior and $\bar{\theta}_{\star}=1.5$ for the uniform one.
@@ E.2 Bayesian logistic regression
Algorithm Approx./Iterative NMSE (%) Times (s)
Laplace Unif Laplace Unif
PPGD Approx $14.70\pm 4.42$ $3.63\pm 4.93$ $102.6\pm 5.1$ $107.9\pm 5.5$
Iterative $19.04\pm 1.34$- $122.3\pm 5.1$-
PIPULA Approx $12.18\pm 1.62$ $4.71\pm 6.02$ $98.8\pm 5.7$ $101.0\pm 4.0$
Iterative $19.22\pm 1.28$- $126.2\pm 3.8$-
Iterative $4.44\pm 1.40$- $129.7\pm 15.8$-
MYIPLA Approx $4.42\pm 1.32$ $15.26\pm 4.44$ ${89.9\pm 4.2}$ ${97.0\pm 4.2}$
Iterative $4.67\pm 1.60$- $120.5\pm 10.1$-
PIPGLA Approx $2.30\pm 0.58$ $6.83\pm 3.97$ $116.5\pm 5.5$ $103.1\pm 8.0$
Iterative $\mathbf{2.02\pm 0.54}$- $122.9\pm 6.9$-
IPLA $7.76\pm 3.39$ $20.12\pm 2.88$ $\mathbf{81.1\pm 3.0}$ $\mathbf{82.9\pm 4.9}$
Results.
Table 6 extends the results in Table 2 by also including the results for PPGD, PIPULA and IPLA (as a benchmark). Figure 4 shows the $\theta$-iterates obtained with MYIPLA and PIPGLA starting from 7 different initial values $\theta_{0}$ and using the approximate solver for $\operatorname{prox}_{g_{2}}^{\lambda}$ with $g_{2}(\theta,x)=\sum_{i=1}^{d_{x}}|x_{i}-\theta|$ and an iterative procedure using 40 iterations in each step. We observe that the iterative solver results in a slightly slower convergence to stationarity, but overall the two sets of algorithms converge to the same true value of $\theta$. We also observe that the convergence to stationarity for PIPGLA is much slower compared to MYIPLA. However, if we increase the value of $\gamma$ in the hope of faster convergence, the iterates either do not converge to the true value or the standard deviation is significantly larger. For all algorithms considered, approximate solvers are 25% faster than iterative solvers (see Table 6).
We also compare the results for the uniform prior, in this case we only use the approximate proximity map (Figure 5), as the iterative approach is not numerically stable.
@@ Results.
(c) PIPGLA, approximated $\operatorname{prox}_{g_{2}}^{\lambda}$.
(d) PIPGLA, iterative solver for $\operatorname{prox}_{g_{2}}^{\lambda}$.
@@
Since all the algorithms considered aim at estimating the parameter $\theta$ by sampling from a distribution which concentrates around $\bar{\theta}_{\star}$, we compare the estimators of $\bar{\theta}_{\star}$ obtained by using only the last iterate $\theta_{K+1}^{N}$ and averaging over a number of iterates. We compare the normalised MSE (NMSE) for $\theta$ for the estimator obtained by averaging the $\theta$-iterates after discarding a burn-in of 1500 samples (column named avg) against using the last $\theta$ of the chain (column last). The results are in agreement, with the NMSE for the averaged estimator having lower variance in most settings (Table 7).
Table 7: Bayesian logistic regression for Laplace and uniform priors. Normalised MSE (NMSE) for the last iterate of $\theta$ (last) and the posterior mean after discarding a burn-in of 1500 samples (avg). Each different algorithm is run 500 times for different starting points using 50 particles and 5000 steps. NMSEs are averaged over the 500 replicates. The second column indicates whether the proximal map is calculated approximately or iteratively, using 40 steps in each iteration. For the uniform prior case, we did not implement the iterative method due to numerical instabilities.
@@
E.3 Bayesian neural network
E.3.1 Sparsity Inducing Prior: MNIST
Our setting is equivalent to assuming that the datapoints’ labels $l$ are conditionally independent given the features $f$ and network weights $x=(w,v)$, and therefore have the following probability density
$p_{\theta}(x,\mathcal{Y}_{\text{train}})=\prod_{i}\text{Laplace}(w_{i}|0,e^{2\alpha})\prod_{j}\text{Laplace}(v_{j}|0,e^{2\beta})\prod_{(f,l)\in\mathcal{Y}_{\text{train}}}p(l|f,x),$
where $x$ denotes the weight matrices, i.e. $x=(w,v)$. We note that the log density can be decomposed as
$-\log p_{\theta}(x,\mathcal{Y}_{\text{train}})=\underbrace{d_{w}\alpha+\sum_{i}|w_{i}|e^{-2\alpha}+d_{v}\beta+\sum_{j}|v_{j}|e^{-2\beta}}_{g_{2}(\theta,x)}\underbrace{-\sum_{(f,l)\in\mathcal{Y}_{\text{train}}}\log p(l|f,x)}_{g_{1}(\theta,x)},$
where $d_{w}$ and $d_{v}$ denote the dimensions of the weights $w$ and $v$, respectively, $g_{1}$ is differentiable and does not depend on $\theta$ and $g_{2}$ is proper, convex and lower semi-continuous. We have derived an approximation to the proximity map of $g_{2}$ in E.1.2.
@@ E.3.1 Sparsity Inducing Prior: MNIST
We use the MNIST dataset. Features are normalised so that each pixel has mean zero and unit standard deviation across the dataset. We split the dataset into 80/20 training and test sets.
@@
$\operatorname{prox}_{g_{2}}^{\lambda}(\alpha,w)=\operatorname*{arg\,min}_{(u_{0},u)}h(u_{0},u),\quad h(u_{0},u)=u_{0}d_{w}+\sum_{i}|u_{i}|e^{-2u_{0}}+\|(u_{0},u)-(\alpha,w)\|^{2}/{(2\lambda)},$
whose approximate solution is calculated in Section E.1.2.
@@
For the $x$-gradients of $g_{1}$, we use JAX’s grad function (implementing a version of autograd). Plugging the expressions above in the corresponding equations, we can implement the proposed algorithms. However, due to the high dimensionality of the latent variables, we stabilise the algorithm using the heuristics discussed in Section 2 of Kuntz et al. [2023]. This simply entails dividing the gradients and proximal mapping terms of the updates of $\alpha$ and $\beta$ by $d_{w}$ and $d_{v}$. We then set $\gamma=0.05$ and $\lambda=0.5$ (after performing a grid search) which ensures that the algorithms are not close to losing stability. In addition, the weights of the network are initialised according to the assumed prior. This is done by setting each weight to ±$\pm\;a\log u$ where $u\sim\mathcal{U}(0,1)$, the sign is chosen uniformly at random and $a>0$ is interpreted as the average initial size of the weights. Williams [1995] suggests setting $a=1/\sqrt{2m}$ for $w$ and $a=1.6/\sqrt{2m}$ for $v$, where $m$ is the fan-in of the destination unit.
@@
and the log pointwise predictive density (LPPD, Vehtari et al. [2017])
$\text{LPPD}:=\frac{1}{|\mathcal{Y}_{\text{test}}|}\sum_{(f,l)\in\mathcal{Y}_{\text{test}}}\log(g(l|f)).$
Under the assumption that data is drawn independently from $p(l,f)$, we have the following approximation for large test data sets,
LPPD $\approx\int\log(g(l|f))p(\mathrm{d}l,\mathrm{d}f)=\int\bigg{[}\int\log\Big{(}\frac{g(l|f)}{p(l|f)}\Big{)}p(\mathrm{d}l|\mathrm{d}f)\bigg{]}p(\mathrm{d}f)+\int\log(p(l|f))p(\mathrm{d}l,\mathrm{d}f)$
@@
This means that the larger the LPPD is, the smaller the mean KL divergence between our classifier $g(l|f)$ and the optimal classifier $p(l|f)$.